# Targeted collapse regularized autoencoder for anomaly detection: black hole at the center

### v1

@@ 4.2 Results
Tables 12, and 3 show anomaly detection results on MNIST, CIFAR-10, and Arrhythmia datasets. On all datasets, our approach achieves the highest performance with a considerable margin. On MNIST and CIFAR-10, it also achieves the highest performance for respectively ten and five out of ten total classes. As the tables show, norm regularization is a simple yet powerful way to improve anomaly detection performance.
@@
Normal OC-SVM scholkopf1999support OCGAN perera2019ocgan VAE kingma2013auto DCAE makhzani2015winner IF liu2008isolation AnoGAN schlegl2017unsupervised KDE parzen1962estimation DSVDD ruff2018deep MEMAE gong2019memorizing MEMGAN yang2021memory Toll 0 98.6$\pm$0.0 99.8 99.7 97.6$\pm$0.7 98.0$\pm$0.3 96.6$\pm$1.3 97.1$\pm$0.0 98.0$\pm$0.7 99.3$\pm$0.1 99.3$\pm$0.1 99.8$\pm$0.1 1 99.5$\pm$0.0 99.9 99.9 98.3$\pm$0.6 97.3$\pm$0.4 99.2$\pm$0.6 98.9$\pm$0.0 99.7$\pm$0.1 99.8$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 2 82.5$\pm$0.1 94.2 93.6 85.4$\pm$2.4 88.6$\pm$0.5 85.0$\pm$2.9 79.0$\pm$0.0 91.7$\pm$0.8 90.6$\pm$0.8 94.5$\pm$0.1 96.2$\pm$1.6 3 88.1$\pm$0.0 96.3 95.9 86.7$\pm$0.9 89.9$\pm$0.4 88.7$\pm$2.1 86.2$\pm$0.0 91.9$\pm$1.5 94.7$\pm$0.6 95.7$\pm$0.4 97.9$\pm$0.6 4 94.9$\pm$0.0 97.5 97.3 86.5$\pm$2.0 92.7$\pm$0.6 89.4$\pm$1.3 87.9$\pm$0.0 94.9$\pm$0.8 94.5$\pm$0.4 96.1$\pm$0.4 97.8$\pm$0.4 5 77.1$\pm$0.0 98.0 96.4 78.2$\pm$2.7 85.5$\pm$0.8 88.3$\pm$2.9 73.8$\pm$0.0 88.5$\pm$0.9 95.1$\pm$0.1 93.6$\pm$0.3 98.1$\pm$0.6 6 96.5$\pm$0.0 99.1 99.3 94.6$\pm$0.5 95.6$\pm$0.3 94.7$\pm$2.7 87.6$\pm$0.0 98.3$\pm$0.5 98.4$\pm$0.5 98.6$\pm$0.1 99.5$\pm$0.1 7 93.7$\pm$0.0 98.1 97.6 92.3$\pm$1.0 92.0$\pm$0.4 93.5$\pm$1.8 91.4$\pm$0.0 94.6$\pm$0.9 95.4$\pm$0.2 96.2$\pm$0.2 98.7$\pm$0.2 8 88.9$\pm$0.0 93.9 92.3 86.5$\pm$1.6 89.9$\pm$0.4 84.9$\pm$2.1 79.2$\pm$0.0 93.9$\pm$1.6 86.9$\pm$0.5 93.5$\pm$0.1 97.3$\pm$0.5 9 93.1$\pm$0.0 98.1 97.6 90.4$\pm$1.8 93.5$\pm$0.3 92.4$\pm$1.1 88.2$\pm$0.0 96.5$\pm$0.3 97.3$\pm$0.2 95.9$\pm$0.1 98.5$\pm$0.3 Average 91.3 97.5 97.0 89.7 92.3 91.4 87.0 94.8 95.2 96.5 98.4
@@
Normal OCGAN VAE DSVDD DSEBM zhai2016deep DAGMM zong2018deep IF AnoGAN ALAD zenati2018adversarially MEMAE MEMGAN Toll airplane 75.7 70.0 61.7$\pm$4.1 41.4$\pm$2.3 56.0$\pm$6.9 60.1$\pm$0.7 67.1$\pm$2.5 64.7$\pm$2.6 66.5$\pm$0.9 73.0$\pm$0.8 75.0$\pm$4.3 auto. 53.1 38.6 65.9$\pm$2.1 57.1$\pm$2.0 48.3$\pm$1.8 50.8$\pm$0.6 54.7$\pm$3.4 38.7$\pm$0.8 46.4$\pm$0.1 52.5$\pm$0.7 67.6$\pm$3.2 bird 64.0 67.9 50.8$\pm$0.8 61.9$\pm$0.1 53.8$\pm$4.0 49.2$\pm$0.4 52.9$\pm$3.0 67.0$\pm$0.7 66.0$\pm$0.1 67.2$\pm$0.1 56.2$\pm$2.8 cat 62.0 53.5 59.1$\pm$1.4 50.1$\pm$0.4 51.2$\pm$0.8 55.1$\pm$0.4 54.5$\pm$1.9 59.2$\pm$0.3 52.9$\pm$0.1 57.3$\pm$0.2 64.6$\pm$1.6 deer 72.3 74.8 60.9$\pm$1.1 73.3$\pm$0.2 52.2$\pm$7.3 49.8$\pm$0.4 65.1$\pm$3.2 72.7$\pm$0.6 72.8$\pm$0.1 73.9$\pm$0.9 67.7$\pm$1.4 dog 62.0 52.3 65.7$\pm$0.8 60.5$\pm$0.3 49.3$\pm$3.6 58.5$\pm$0.4 60.3$\pm$2.6 52.8$\pm$1.2 52.9$\pm$0.2 65.0$\pm$0.2 68.1$\pm$1.6 frog 72.3 68.7 67.7$\pm$2.6 68.4$\pm$0.3 64.9$\pm$1.7 42.9$\pm$0.6 58.5$\pm$1.4 69.5$\pm$1.1 63.7$\pm$0.4 72.8$\pm$0.7 74.7$\pm$2.5 horse 57.5 49.3 67.3$\pm$0.9 53.3$\pm$0.7 55.3$\pm$0.8 55.1$\pm$0.7 62.5$\pm$0.8 44.8$\pm$0.4 45.9$\pm$0.1 52.5$\pm$0.5 65.8$\pm$1.3 ship 82.0 69.6 75.9$\pm$1.2 73.9$\pm$0.3 51.9$\pm$2.4 74.2$\pm$0.6 75.8$\pm$4.1 73.4$\pm$0.4 70.1$\pm$0.1 74.4$\pm$0.3 80.0$\pm$1.1 truck 55.4 38.6 73.1$\pm$1.2 63.6$\pm$3.1 54.2$\pm$5.8 58.9$\pm$0.7 66.5$\pm$2.8 39.2$\pm$1.3 48.2$\pm$0.2 65.6$\pm$1.6 76.2$\pm$0.5 Average 65.6 58.3 64.8 60.4 54.4 55.5 61.8 59.3 58.5 65.3 69.6
[…]
@@
IF OC-SVM DSEBM-e AnoGAN DAGMM ALAD RCGAN yang2020regularized DSVDD GOAD bergman2020classification MEMAE MEMGAN Toll 53.03 45.18 46.01 42.42 49.83 51.52 54.14 34.79 52.00 51.13 55.72 60.00
@@ 4.3 Ablation Study
Table 4 compares the anomaly detection performance with and without norm minimization. The effect of regularization varies across different normal classes. This is due to the fact that different classes benefit to different extents from the regularization and that the single $\beta$ value used across various classes is not necessarily a good fit for all. Even with this caveat, the contribution of norm minimization is significant at about 4 percent on average in this high-accuracy regime. For the given $\beta$, this effect is the most pronounced when considering digit “8” as the normal class. Figure 4 shows the effect of introducing the regularization at various intensities on performance.
@@
Normal $\beta$=0 $\beta$=1000
0 99.1$\pm$0.1 99.8$\pm$0.1
1 99.9$\pm$0.0 99.9$\pm$0.0
2 91.4$\pm$0.3 96.2$\pm$1.6
3 93.4$\pm$0.2 97.9$\pm$0.6
4 94.3$\pm$0.3 97.8$\pm$0.4
5 95.1$\pm$0.1 98.1$\pm$0.6
6 98.5$\pm$0.1 99.5$\pm$0.1
7 96.4$\pm$0.1 98.7$\pm$0.2
8 82.9$\pm$0.2 97.3$\pm$0.5
9 96.2$\pm$0.1 98.5$\pm$0.3
Average 94.7 98.4
@@ 5 Conclusion
In this work, we proposed a simple method for regularizing autoencoder training and modifying the anomaly score function based on minimizing the norm of bottleneck representations. The method comes from simple first principles and despite its simplicity, appropriately fits the purpose of anomaly detection, leading to a significant performance boost. The implementation is straightforward and the methodology does not introduce any constraints on the application or the associated data modality. We showcased the technique’s advantages by conducting comparisons of our method against well-established and intricate approaches to anomaly detection, as well as through ablation studies that evaluate the impact of the regularization effect. We also carried out an analytical investigation into the learning dynamics, exploring the phenomenological consequences of the proposed modification and how it complements the capabilities of autoencoders.

### v2

@@ 4.2 Results
Tables 123, and 4 show anomaly detection results on the datasets. On all datasets, our approach achieves a competitive performance and outperforms several well-known and considerably more sophisticated methods in various normal class experiments and on average for some datasets. Notably, GeoTrans [23] and ARNet [30] incorporate significantly larger, more sophisticated network architectures, as well as data augmentation and self-supervised training objectives. As the tables show, norm regularization is a simple yet powerful way to improve anomaly detection performance.
Tables 5 and 6 demonstrate the effect of targeted collapse on the performance of FITYMI as a strong anomaly detection technique based on using powerful pre-trained backbones and synthetic out-of-distribution data. Even in this high-performance regime, the addition of targeted collapse is capable of further improving the model performance virtually at no cost and despite the availability of out-of-distribution data that can already largely inform the model of the desirable separation in the latent space.
@@
Normal OC-SVM [56] OCGAN [48] VAE [33] DCAE [40] IF [37] AnoGAN [55] KDE [47] DSVDD [53] MEMAE [24] MEMGAN [66] Toll 0 98.6$\pm$0.0 99.8 99.7 97.6$\pm$0.7 98.0$\pm$0.3 96.6$\pm$1.3 97.1$\pm$0.0 98.0$\pm$0.7 99.3$\pm$0.1 99.3$\pm$0.1 99.8$\pm$0.1 1 99.5$\pm$0.0 99.9 99.9 98.3$\pm$0.6 97.3$\pm$0.4 99.2$\pm$0.6 98.9$\pm$0.0 99.7$\pm$0.1 99.8$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 2 82.5$\pm$0.1 94.2 93.6 85.4$\pm$2.4 88.6$\pm$0.5 85.0$\pm$2.9 79.0$\pm$0.0 91.7$\pm$0.8 90.6$\pm$0.8 94.5$\pm$0.1 96.2$\pm$1.6 3 88.1$\pm$0.0 96.3 95.9 86.7$\pm$0.9 89.9$\pm$0.4 88.7$\pm$2.1 86.2$\pm$0.0 91.9$\pm$1.5 94.7$\pm$0.6 95.7$\pm$0.4 97.9$\pm$0.6 4 94.9$\pm$0.0 97.5 97.3 86.5$\pm$2.0 92.7$\pm$0.6 89.4$\pm$1.3 87.9$\pm$0.0 94.9$\pm$0.8 94.5$\pm$0.4 96.1$\pm$0.4 97.8$\pm$0.4 5 77.1$\pm$0.0 98.0 96.4 78.2$\pm$2.7 85.5$\pm$0.8 88.3$\pm$2.9 73.8$\pm$0.0 88.5$\pm$0.9 95.1$\pm$0.1 93.6$\pm$0.3 98.1$\pm$0.6 6 96.5$\pm$0.0 99.1 99.3 94.6$\pm$0.5 95.6$\pm$0.3 94.7$\pm$2.7 87.6$\pm$0.0 98.3$\pm$0.5 98.4$\pm$0.5 98.6$\pm$0.1 99.5$\pm$0.1 7 93.7$\pm$0.0 98.1 97.6 92.3$\pm$1.0 92.0$\pm$0.4 93.5$\pm$1.8 91.4$\pm$0.0 94.6$\pm$0.9 95.4$\pm$0.2 96.2$\pm$0.2 98.7$\pm$0.2 8 88.9$\pm$0.0 93.9 92.3 86.5$\pm$1.6 89.9$\pm$0.4 84.9$\pm$2.1 79.2$\pm$0.0 93.9$\pm$1.6 86.9$\pm$0.5 93.5$\pm$0.1 97.3$\pm$0.5 9 93.1$\pm$0.0 98.1 97.6 90.4$\pm$1.8 93.5$\pm$0.3 92.4$\pm$1.1 88.2$\pm$0.0 96.5$\pm$0.3 97.3$\pm$0.2 95.9$\pm$0.1 98.5$\pm$0.3 Average 91.3 97.5 97.0 89.7 92.3 91.4 87.0 94.8 95.2 96.5 98.4
@@
Normal OCGAN VAE DSVDD DSEBM [68] DAGMM [73] IF AnoGAN ALAD [67] MEMAE MEMGAN Toll airplane 75.7 70.0 61.7$\pm$4.1 41.4$\pm$2.3 56.0$\pm$6.9 60.1$\pm$0.7 67.1$\pm$2.5 64.7$\pm$2.6 66.5$\pm$0.9 73.0$\pm$0.8 75.0$\pm$4.3 auto. 53.1 38.6 65.9$\pm$2.1 57.1$\pm$2.0 48.3$\pm$1.8 50.8$\pm$0.6 54.7$\pm$3.4 38.7$\pm$0.8 46.4$\pm$0.1 52.5$\pm$0.7 67.6$\pm$3.2 bird 64.0 67.9 50.8$\pm$0.8 61.9$\pm$0.1 53.8$\pm$4.0 49.2$\pm$0.4 52.9$\pm$3.0 67.0$\pm$0.7 66.0$\pm$0.1 67.2$\pm$0.1 56.2$\pm$2.8 cat 62.0 53.5 59.1$\pm$1.4 50.1$\pm$0.4 51.2$\pm$0.8 55.1$\pm$0.4 54.5$\pm$1.9 59.2$\pm$0.3 52.9$\pm$0.1 57.3$\pm$0.2 64.6$\pm$1.6 deer 72.3 74.8 60.9$\pm$1.1 73.3$\pm$0.2 52.2$\pm$7.3 49.8$\pm$0.4 65.1$\pm$3.2 72.7$\pm$0.6 72.8$\pm$0.1 73.9$\pm$0.9 67.7$\pm$1.4 dog 62.0 52.3 65.7$\pm$0.8 60.5$\pm$0.3 49.3$\pm$3.6 58.5$\pm$0.4 60.3$\pm$2.6 52.8$\pm$1.2 52.9$\pm$0.2 65.0$\pm$0.2 68.1$\pm$1.6 frog 72.3 68.7 67.7$\pm$2.6 68.4$\pm$0.3 64.9$\pm$1.7 42.9$\pm$0.6 58.5$\pm$1.4 69.5$\pm$1.1 63.7$\pm$0.4 72.8$\pm$0.7 74.7$\pm$2.5 horse 57.5 49.3 67.3$\pm$0.9 53.3$\pm$0.7 55.3$\pm$0.8 55.1$\pm$0.7 62.5$\pm$0.8 44.8$\pm$0.4 45.9$\pm$0.1 52.5$\pm$0.5 65.8$\pm$1.3 ship 82.0 69.6 75.9$\pm$1.2 73.9$\pm$0.3 51.9$\pm$2.4 74.2$\pm$0.6 75.8$\pm$4.1 73.4$\pm$0.4 70.1$\pm$0.1 74.4$\pm$0.3 80.0$\pm$1.1 truck 55.4 38.6 73.1$\pm$1.2 63.6$\pm$3.1 54.2$\pm$5.8 58.9$\pm$0.7 66.5$\pm$2.8 39.2$\pm$1.3 48.2$\pm$0.2 65.6$\pm$1.6 76.2$\pm$0.5 Average 65.6 58.3 64.8 60.4 54.4 55.5 61.8 59.3 58.5 65.3 69.6
Table 3: Mean and standard deviation of AUC in % on Fashion-MNIST and CIFAR-100 classes over 10 seeds. The highest performance in each row is in bold.
Dataset Normal DAGMM DSEBM ADGAN [12] GANomaly [2] GeoTrans [23] ARNet [30] Toll T-shirt/top 42.1 91.6 89.9 80.3 99.4 92.7 91.9$\pm$0.2 Trouser 55.1 71.8 81.9 83.0 97.6 99.3 98.7$\pm$0.0 Pullover 50.4 88.3 87.6 75.9 91.1 89.1 91.3$\pm$0.2 Dress 57.0 87.3 91.2 87.2 89.9 93.6 92.4$\pm$0.2 Fashion-MNIST Coat 26.9 85.2 86.5 71.4 92.1 90.8 91.5$\pm$0.1 Sandal 70.5 87.1 89.6 92.7 93.4 93.1 93.1$\pm$0.2 Shirt 48.3 73.4 74.3 81.0 83.3 85.0 84.6$\pm$0.3 Sneaker 83.5 98.1 97.2 88.3 98.9 98.4 99.0$\pm$0.0 Bag 49.9 86.0 89.0 69.3 90.8 97.8 90.8$\pm$0.6 Ankle boot 34.0 97.1 97.1 80.3 99.2 98.4 98.9$\pm$0.1 Average 51.8 86.6 88.4 80.9 93.5 93.9 93.2 aquatic mammals 43.4 64.0 63.1 57.9 74.7 77.5 70.3$\pm$3.4 fish 49.5 47.9 54.9 51.9 68.5 70.0 66.2$\pm$1.1 flowers 66.1 53.7 41.3 36.0 74.0 62.4 80.3$\pm$0.7 food containers 52.6 48.4 50.0 46.5 81.0 76.2 58.2$\pm$1.5 fruit and vegetables 56.9 59.7 40.6 46.6 78.4 77.7 80.1$\pm$1.3 household electrical devices 52.4 46.6 42.8 42.9 59.1 64.0 61.4$\pm$2.6 household furniture 55.0 51.7 51.1 53.7 81.8 86.9 64.8$\pm$3.0 insects 52.8 54.8 55.4 59.4 65.0 65.6 70.5$\pm$1.3 large carnivores 53.2 66.7 59.2 63.7 85.5 82.7 62.1$\pm$2.1 CIFAR-100 large man-made outdoor things 42.5 71.2 62.7 68.0 90.6 90.2 79.6$\pm$1.4 large natural outdoor scenes 52.7 78.3 79.8 75.6 87.6 85.9 81.6$\pm$1.1 large omnivores and herbivores 46.4 62.7 53.7 57.6 83.9 83.5 62.6$\pm$1.2 medium-sized mammals 42.7 66.8 58.9 58.7 83.2 84.6 61.2$\pm$1.9 non-insect invertebrates 45.4 52.6 57.4 59.9 58.0 67.6 58.6$\pm$2.0 people 57.2 44.0 39.4 43.9 92.1 84.2 65.4$\pm$4.5 reptiles 48.8 56.8 55.6 59.9 68.3 74.1 56.1$\pm$1.5 small mammals 54.4 63.1 63.3 64.4 73.5 80.3 62.3$\pm$1.1 trees 36.4 73.0 66.7 71.8 93.8 91.0 80.9$\pm$1.8 vehicles 1 52.4 57.7 44.3 54.9 90.7 85.3 65.2$\pm$1.4 vehicles 2 50.3 55.5 53.0 56.8 85.0 85.4 63.6$\pm$1.6 Average 50.5 58.8 54.7 56.5 78.7 78.8 67.6
Table 4: Average F1 score in % on Arrhythmia dataset over 10 seeds. The highest performance is in bold.
@@
IF OC-SVM DSEBM-e AnoGAN DAGMM ALAD RCGAN [65] DSVDD GOAD [7] MEMAE MEMGAN Toll 53.03 45.18 46.01 42.42 49.83 51.52 54.14 34.79 52.00 51.13 55.72 60.00
Table 5: The effect of targeted collapse on the performance of FITYMI evaluated on CIFAR-10. Reported results are the mean and standard deviation of AUC in % for each class over 10 seeds. The highest performance in each experiment is in bold.
Normal 0 1 2 3 4 5 6 7 8 9 Average FITYMI [41] 99.25$\pm$0.07 98.63$\pm$0.26 97.08$\pm$0.94 98.10$\pm$0.17 98.93$\pm$0.27 96.06$\pm$0.56 99.36$\pm$0.10 99.53$\pm$0.05 98.88$\pm$0.55 98.29$\pm$1.23 98.41$\pm$0.19 FITYMI+Toll 99.17$\pm$0.05 99.23$\pm$0.21 97.42$\pm$0.38 98.14$\pm$0.17 99.23$\pm$0.10 97.25$\pm$0.42 99.60$\pm$0.05 99.64$\pm$0.06 99.48$\pm$0.14 98.62$\pm$0.17 98.78$\pm$0.07
Table 6: The effect of targeted collapse on the performance of FITYMI evaluated on CIFAR-100. Reported results are the mean and standard deviation of AUC in % for each class over 10 seeds. The highest performance in each experiment is in bold.
Normal 0 1 2 3 4 5 6 7 8 9 10 FITYMI [41] 95.41$\pm$0.30 95.44$\pm$0.80 97.56$\pm$0.75 96.02$\pm$1.05 96.51$\pm$0.34 96.09$\pm$0.67 98.42$\pm$1.09 96.02$\pm$0.34 97.36$\pm$0.21 96.93$\pm$0.32 96.44$\pm$0.51 FITYMI+Toll 97.29$\pm$0.32 96.21$\pm$0.94 97.57$\pm$1.67 96.30$\pm$1.50 97.33$\pm$0.22 97.05$\pm$0.44 99.02$\pm$0.06 96.83$\pm$0.51 98.43$\pm$0.11 96.84$\pm$0.65 95.86$\pm$0.70 Normal 11 12 13 14 15 16 17 18 19 Average FITYMI [41] 96.91$\pm$0.21 96.89$\pm$0.20 93.43$\pm$0.62 96.93$\pm$1.41 95.12$\pm$0.50 96.88$\pm$0.25 96.25$\pm$0.73 98.46$\pm$0.29 96.12$\pm$0.44 96.46$\pm$0.11 FITYMI+Toll 97.73$\pm$0.12 97.80$\pm$0.17 93.99$\pm$0.94 97.18$\pm$0.95 96.69$\pm$0.32 97.53$\pm$0.15 96.82$\pm$1.11 98.74$\pm$0.16 97.04$\pm$0.29 97.11$\pm$0.14
@@ 4.3 Ablation Study
Table 7 compares the anomaly detection performance with and without norm minimization. The effect of regularization varies across different normal classes. This is due to the fact that different classes benefit to different extents from the regularization and that the single $\beta$ value used across various classes is not necessarily a good fit for all. Even with this caveat, the contribution of norm minimization is significant at about 4 percent on average in this high-accuracy regime. For the given $\beta$, this effect is the most pronounced when considering digit “8” as the normal class. Figure 4 shows the effect of introducing the regularization on performance at various intensities.
@@
Normal $\beta$=0 $\beta$=1000 0 99.1$\pm$0.1 99.8$\pm$0.1 1 99.9$\pm$0.0 99.9$\pm$0.0 2 91.4$\pm$0.3 96.2$\pm$1.6 3 93.4$\pm$0.2 97.9$\pm$0.6 4 94.3$\pm$0.3 97.8$\pm$0.4 5 95.1$\pm$0.1 98.1$\pm$0.6 6 98.5$\pm$0.1 99.5$\pm$0.1 7 96.4$\pm$0.1 98.7$\pm$0.2 8 82.9$\pm$0.2 97.3$\pm$0.5 9 96.2$\pm$0.1 98.5$\pm$0.3 Average 94.7 98.4
@@
In our method, we have incorporated the Euclidean norm. In principle, other norms such as $L_{1}$ can also be adopted. Table 8 demonstrates the effect of the choice of norm for the MNIST dataset. For the $L_{1}$ and $L_{2}$ norms, we obtain a similar level of performance, while performance is lowered with $L_{\infty}$.
Table 8: AUC comparison for MNIST classes over 10 seeds for using $L_{1}$, $L_{2}$, and $L_{\infty}$ norms of the bottleneck representations. The highest performance in each row is in bold.
Normal $1$-norm $2$-norm $\infty$-norm 0 99.7$\pm$0.1 99.8$\pm$0.1 99.3$\pm$0.3 1 99.8$\pm$0.0 99.9$\pm$0.0 99.7$\pm$0.1 2 96.4$\pm$1.6 96.2$\pm$1.6 94.4$\pm$2.0 3 97.8$\pm$0.8 97.9$\pm$0.6 97.3$\pm$0.4 4 97.8$\pm$0.3 97.8$\pm$0.4 96.8$\pm$0.2 5 98.3$\pm$0.6 98.1$\pm$0.6 96.9$\pm$0.8 6 99.5$\pm$0.1 99.5$\pm$0.1 99.1$\pm$0.3 7 98.6$\pm$0.3 98.7$\pm$0.2 97.6$\pm$0.3 8 97.1$\pm$0.5 97.3$\pm$0.5 95.2$\pm$0.8 9 98.5$\pm$0.3 98.5$\pm$0.3 97.3$\pm$0.3 Average 98.4 98.4 97.4
@@ 5 Conclusion
In this work, we proposed a simple method for regularizing autoencoder training and modifying the anomaly score function based on minimizing the norm of bottleneck representations. The method comes from simple first principles and despite its simplicity, appropriately fits the purpose of anomaly detection, leading to a significant performance boost. The implementation is straightforward and the methodology does not introduce any constraints on the application or the associated data modality. We showcased the technique’s advantages by conducting comparisons of our method against well-established and intricate approaches to anomaly detection, as well as through ablation studies that evaluate the impact of the regularization effect. We further demonstrated that the idea of targeted collapse can sit well with cutting-edge techniques and can still improve their performance despite their extensively pre-trained foundation models and access to anomalous samples during training. We also carried out an analytical investigation into the learning dynamics, exploring the phenomenological consequences of the proposed modification and how it complements the capabilities of autoencoders.